# Investigating MAXI J1752--457 with NuSTAR in the aftermath of a superburst

### v1

@@ Abstract
We present two NuSTAR observations of the X-ray transient, MAXI J1752--457, following a superburst which was observed by MAXI/GSC in November, 2024. NuSTAR follow-up confirmed that MAXI J1752--457 is coincident with the previously observed Einstein Probe source, EP240809a. We performed a spectral analysis of the source during both NuSTAR observations, and we confirm that the hard X-ray spectra are consistent with the inclusion of a spherical blackbody component. At about 79 hours after the onset of the superburst, we find a blackbody temperature of $kT_{\mathrm{bb}}=0.60\pm 0.1$ keV and $R_{\mathrm{bb}}/D_{8}=6.0^{+0.4}_{-0.3}$ km (not including corrections for scattering in the neutron star atmosphere), where $D_{8}$ is the source distance, which is not yet known, in units of 8 kpc. Furthermore, we report a hard X-ray excess which is fit well by a power law with photon index $\Gamma\approx 4$, much steeper than those typically seen during accretion onto neutron stars at similar luminosities. We infer that the electron energy distribution in the Comptonizing medium which gives rise to the power law component differs significantly compared to purely accretion-powered neutron star transients even several days after the onset and rapid decay of superbursts. We also performed an energy-resolved timing analysis which showed that the source variability was dominated by red noise in the Comptonization component, suggesting coupling with an accretion disk, rather than being seeded by thermal emission from the neutron star surface.
@@
facilities: MAXI/GSC, NuSTARsoftware: astropy (Astropy Collaboration et al., 2013, 2018, 2022), Stingray (Huppenkothen et al., 2019; Bachetti et al., 2024a, b), Xspec (Arnaud, 1996), SAOImageDS9 (Joye & Mandel, 2003)
1 Introduction
@@
X-ray binaries (XRB), formed via accretion onto a compact object from a companion, exhibit a wide array of variability on timescales ranging from milliseconds to days to months. In particular, most XRBs are transient, meaning that they alternate between intervals of quiescence and outburst due to changes in the rate of mass accretion. For example, in Be X-ray binaries consisting of a Be star and an accreting neutron star, these periods of outburst are caused by transient accretion when the compact object passes near its companion in its orbit. XRBs hosting black holes undergo weeks- or months-long outbursts due to slow accumulation of material into their accretion disks, leading to a critical point at which material accretes much more efficiently onto the compact object (Mineshige & Wheeler, 1989).
@@
These accretion-powered outbursts are distinct from bursting behavior caused by thermonuclear burning. Neutron star XRBs which persistently accrete from low mass companions at a relatively low level may accumulate material on their surfaces without burning until a critical depth is reached, producing runaway burning of accreted hydrogen and helium. We observe this runaway conflagration in the form of a rapid flash known as a Type I X-ray burst, lasting only seconds to minutes, during which the source luminosity may increase by multiple orders of magnitude up to the Eddington luminosity. The depth at which accumulated material ignites plays a decisive role in the duration and structure of the resulting burst. Some sources exhibit “superbursts” which resemble Type I X-ray bursts but which last several hours and decay on the timescale of days due to ignition at much larger column depths producing extended carbon burning (in’t Zand, 2017; Cumming, 2003; Keek et al., 2008; Serino et al., 2016). Due to their extended timescale, superbursts can provide an opportunity to probe how accretion disks and coronae respond to high levels of X-ray irradiation. The rarity of these events, however, has thus far limited investigations into superbursts. In this paper, we present an analysis of the newly-discovered source, MAXI J1752--457, which was observed by the Nuclear Spectroscopic Telescope Array (NuSTAR) during the decay phase of a superburst.
@@
The first detection of MAXI J1752--457 with the Monitor of All-sky X-ray Image (MAXI) Gas Slit Camera (GSC) occurred during the scan transit at 18:23 on November 9, 2024 UTC (MJD 60623.766) with a 4–10 keV flux of $982\pm 50$ mCrab, or $(1.12\pm 0.06)\times 10^{-8}$$\mathrm{erg\,cm^{-2}\,s^{-1}}$(Serino et al., 2024). The flux dropped rapidly over subsequent scan transits. Several months prior to the detection of MAXI J1752--457 with MAXI, an X-ray transient, deemed EP240809a, was detected by Einstein Probe (EP) with a position consistent with MAXI J1752--457 (Liu et al., 2024). This earlier period of brightening, which was observed on August 9, 2024 (UTC), reached a much lower flux than the November burst: Liu et al. (2024) reported an unabsorbed 0.4–5 keV flux of $5.3\times 10^{-11}$$\mathrm{erg\,cm^{-2}\,s^{-1}}$, corresponding to about 1.5 mCrab. An optical flare from the source was also observed by the Asteroid Terrestrial-impact Last Alert System (ATLAS) beginning about three weeks before the EP detection (Zhu et al., 2024). The authors pointed out that the soft X-ray and the blue optical spectra were consistent with a tidal disruption event (TDE), but no host galaxy was found. Therefore, the nature of the source remained unknown.
@@
MAXI/GSC observations showed a thermal spectrum which softened during the rapid decay, leading Negoro et al. (2024) to propose that the origin of the the source activity was a long X-ray burst from an uncatalogued X-ray binary, rather than a TDE. By combining MAXI/GSC and NinjaSat observations of the November burst, Aoyama et al. (2025) have shown that the source is indeed consistent with an accreting neutron star. The flux evolution of the burst, which exhibits a fast rise and exponential decay on the timescale of days, along with an X-ray spectrum which can be described with a spherical blackbody, indicate that the transient event corresponded to accretion-induced thermonuclear burning. In particular, the hours-long timescale of the burst and the years-long recurrence time (inferred from the lack of previous detections of the source by MAXI), indicate a superburst.
@@
At the time that it was detected by MAXI/GSC, MAXI J1752--457 had a small Solar Aspect Angle (SAA) of $<45^{\circ}$, making follow-up observations impossible for most pointing X-ray observatories. Thanks to the ability of NuSTAR to point near or at the Sun, we were able to trigger hard X-ray observations of the MAXI confidence region. These observations confirmed that MAXI J1752--457 is coincident with EP240809a, and preliminary spectral analysis provided the first evidence for a hard X-ray excess in addition to the blackbody spectral component (Pike et al., 2024).
@@
Here, we present a detailed analysis of two NuSTAR observations of MAXI J1752--457 following its November, 2024, superburst, and we place the flux of the source as measured by NuSTAR in the context of MAXI/GSC monitoring of the burst. In Section 2, we briefly present the MAXI/GSC observations of the burst. In Section 3, we present both spectral and timing analyses of the source with NuSTAR. Finally, in Section 4, we discuss our results and conclusions. All errors are quoted at the $90\%$ confidence level unless otherwise stated.
2 MAXI light-curve
@@
MAXI, launched in 2009, is an all-sky X-ray monitor installed on the Japanese Experiment Module onboard the International Space Station (Matsuoka et al., 2009). It consists of two instruments: the Gas Slit Camera (GSC) with a net bandpass of 2–20 keV and field of view (FOV) of $1.5^{\circ}\times 160^{\circ}$ (Mihara et al., 2011), and the Solid-state Slit Camera (SSC) with a net bandpass of 0.7–7 keV and FOV of $1.5^{\circ}\times 90^{\circ}$. The SSC ceased operations on June 22, 2022, and was therefore not used during observations of MAXI J1752--457. MAXI/GSC scans the full X-ray sky once every orbit (about 92 minutes), such that individual sources are within the FOV for about $45$ seconds per orbit (Sugizaki et al., 2011). The resulting 3-$\sigma$ detection limit for one scan is $80$ mCrab in the 3–10 keV band (Negoro et al., 2016). The MAXI/GSC light-curve of MAXI J1752--457 available from the online MAXI light-curve portal [^1] http://maxi.riken.jp/star_data/J1752-457/J1752-457.html contained data with a low signal-to-noise ratio due to contamination by the bright low-mass XRB (LMXB), H 1735--444, separated from MAXI J1752--457 by $2.8^{\circ}$ in the GSC anode direction. Therefore, we instead obtained the MAXI/GSC light-curve using the point-spread-function fit method presented by Morii et al. (2016). We show the MAXI/GSC light-curve for MJD 60622 – MJD 60629 in the top panel of Figure 1. We have re-binned the data by up to 10 orbits for clarity before and after the brightest phase of the burst. We note that visual inspection of the MAXI/GSC light-curve of MAXI J1752--457 binned into 16 days over the course of more than 15 years reveals a hint of a weak enhancement ($\sim$ 1–2 mCrab) beginning in July, 2024, and lasting roughly through November, 2024.
3 NuSTAR Data and Analysis
@@
MAXI J1752--457 was observed twice by NuSTAR during the tail of the 2024 superburst. The first observation was performed using a tiling strategy designed to locate the source within the MAXI/GSC 90% confidence region. Four tiling observations were carried out with exposures of about $5$ ks each. The source was detected in one of these observations with OBSID 81010302001. Using the resulting localization, a second, pointed observation (OBSID 81001347001) was carried out less than a day later. The start and end times and the total exposure times of the two NuSTAR observations are listed in Table 1. We extracted cleaned event lists for both observations using the standard nupipeline options with saamode=NONE and tentacle=no. We used HEASOFT v6.34, NuSTARDAS v2.1.4a, and CALDB v20241104.
@@
During both NuSTAR observations, the source was at a low SAA, leading to complications in the image reconstruction process. At low SAA, the primary star tracker, Camera Head Unit 4 (CHU4), which is located on the X-ray optics bench, may be blinded by the Sun and therefore become unusable for image reconstruction. The usual science mode, which utilizes CHU4, is called Mode 1. Outside of Mode 1, image reconstruction must be performed using combinations of CHUs 1, 2, and 3, which are located on the spacecraft bus. This is referred to as Mode 06 data. Depending on the pointing of the spacecraft any given time, one of seven combinations of these three star trackers may be used to produce an attitude solution. Due to thermal flexing of the mountings of CHUs 1, 2, and 3, the resulting solution may include inaccuracies which manifest as a “smearing” of the source or the presence of multiple centroids in the reconstructed image. In order to most effectively extract scientific products from Mode 06 data, we used the nusplitsc command to split the cleaned event list according to active CHU combination. We performed nusplitsc two separate times, once with splitmode=NORMAL and another time with splitmode=STRICT. In some cases, strict splitting can help to remove flickering between CHU combinations, therefore mitigating multiple centroids. However, based on visual inspection using SAOImageDS9 (Joye & Mandel, 2003), we determined that using the strict splitmode did not improve the image reconstruction compared to normal mode, and in fact strict mode only resulted in a reduction in the overall number of events available for science products. We therefore proceeded using the event files produced using splitmode=NORMAL.
@@
We used DS9 to inspect the event files from Mode 1 and the split event files from Mode 06. For each file, we chose a circular source extraction region and centered it on the source using the DS9 centroid calculation feature. During the first observation, there was significant stray light which did not overlap with the source position, but which was close enough to the source to become a consideration when choosing the source region. In order to avoid stray light contamination of the source data, we chose source regions with radius $35^{\prime\prime}$ for FPMA data and $40^{\prime\prime}$ for FPMB data from the first observation. For both FPMA and FPMB data from the second observation, we chose source regions with radius $45^{\prime\prime}$. For all event files, we chose circular background regions with radius $100^{\prime\prime}$. The stray light region, characteristic centroid smearing, and examples of the source and background regions we chose for FPMA are shown in the Mode 06 images for both observations in Figure 2. After selecting extraction regions, we used nuproducts to produce source and background spectra and light-curves for each cleaned event file. The background-subtracted light-curves (summed FPMA and FPMB) are shown in the bottom panels of Figure 1. Finally, for each observation, we used addspec to sum the Mode 1 and Mode 06 data into one pair of FPMA source and background spectra and one pair of FPMB spectra. We used ftgrouppha to rebin the spectra according to the method described by Kaastra & Bleeker (2016) with a minimum of 10 counts per bin.
3.1 Spectral Analysis
@@
Table 2: Spectral parameters and unabsorbed 2–10 keV flux measured during the NuSTAR observations.
@@
$R\,\mathrm{(km)}$$\dagger$\$$\dagger$\$footnotemark: $7.7^{+1.2}_{-1.4}$ $6.0^{+0.4}_{-0.3}$
@@
footnotetext: Radius calculated assuming a distance of $8$ kpc, without correcting for scattering in the neutron star atmosphere.
@@
Given the superburst interpretation presented in Aoyama et al. (2025), we added a physically-motivated single-temperature blackbody component to our spectral models, representing thermal emission from the surface of a neutron star heated by either thermonuclear burning during the burst or by persistent accretion flow onto a boundary layer. Notably, the peak flux observed by MAXI/GSC during the burst was around $2\times 10^{-8}$$\mathrm{erg\,cm^{-2}\,s^{-1}}$. Although the distance to MAXI J1752--457 is unknown, the source is located at $(\alpha,\delta)=(268.269,-45.866)$, or $(l,b)=(346.092,-9.8517)$, placing it at a $17^{\circ}$ separation from the Galactic center. Like Aoyama et al. (2025), we tentatively assume that the angular proximity of the source to the Galactic center translates into physical proximity, giving a distance of around 8 kpc. This assumption leads to a derived peak luminosity on the order of $10^{38}$$\mathrm{erg\,s^{-1}}$, very near the Eddington luminosity for spherical accretion onto a neutron star: $L_{\mathrm{Edd}}=1.38\left(M/{M_{\odot}}\right)\times 10^{38}$$\mathrm{erg\,s^{-1}}$, where the canonical neutron star mass is $M=1.4M_{\odot}$. Thus, the physical interpretation of the source as a bursting neutron star is consistent with a distance of around 8 kpc, and we therefore present derived luminosities and blackbody radii using this fiducial value.
@@
We also considered slightly different forms of this basic spectral model. For example, we replaced the single-temperature blackbody with a multi-temperature disk blackbody. This model resulted in a comparable but slightly worse test statistic. This model yields an inner disk temperature of around $kT_{\mathrm{in}}=0.7$ keV and a slightly shallower power law with a photon index of $\Gamma=3.4$. Similarly, replacing the simple power law component with a thermally Comptonized continuum component (nthcomp; Zdziarski et al., 1996; Życki et al., 1999), did not improve the test statistic. In this case, the relevant spectral parameters did not change compared to the simple power law model.
In order to check the self-consistency of the spectral model, we also reproduced the spectral fits using simpl (Steiner et al., 2009) rather than an analytic power law. We again found no appreciable change in the test statistic. This model yields a slightly larger blackbody radius (at 8 kpc, uncorrected for atmospheric scattering effects) of $11^{+2}_{-1}$ km and $7.4^{+0.6}_{-0.5}$ km for the first and second NuSTAR observations, respectively. For both observations, using simpl instead of an analytic power law results in a decrease in the blackbody temperature by a few percent, an increase in the photon index of about 0.1, and an upscattering fraction of around $10\%$, assuming that the power law is seeded by the spherical blackbody component.

### v2

# Investigating MAXI J1752-457 with NuSTAR in the aftermath of a superburst

@@ Abstract
We present two NuSTAR observations of the X-ray transient, MAXI J1752-457, following a superburst which was observed by MAXI/GSC in November, 2024. NuSTAR follow-up confirmed that MAXI J1752-457 is coincident with the previously observed Einstein Probe source, EP240809a. We performed a spectral analysis of the source during both NuSTAR observations, and we find that the hard X-ray spectra are consistent with the inclusion of a spherical blackbody component and a steep, non-thermal, power law component. At about 79 hours after the onset of the superburst, we find a blackbody temperature of $kT_{\mathrm{bb}}=0.60\pm 0.1$ keV and $R_{\mathrm{bb}}/D_{8}=6.0^{+0.4}_{-0.3}$ km (not including corrections for scattering in the neutron star atmosphere), where $D_{8}$ is the source distance, which is not yet known, in units of 8 kpc. We found that the blackbody temperature did not change significantly in the one day interval between successive NuSTAR observations, and we performed an energy-resolved timing analysis which showed that the source variability was dominated by red noise in the power law component, suggesting coupling with an accretion disk. We infer that the source had entered an accretion-powered flux state. Furthermore, we measure a photon index of $\Gamma\approx 4$, much steeper than those typically observed during accretion onto neutron stars at similar luminosities. This is suggestive of ongoing evolution of the electron energy distribution responsible for the power law component several days after the superburst, but the lack of hard X-ray observations prior to and throughout the superburst make it difficult to present a conclusive physical interpretation of this result.
@@
facilities: MAXI/GSC, NuSTARsoftware: astropy (Astropy Collaboration et al., 2013, 2018, 2022), Stingray (Huppenkothen et al., 2019; Bachetti et al., 2024b, a), Xspec (Arnaud, 1996), SAOImageDS9 (Joye and Mandel, 2003)
I Introduction
@@
X-ray binaries (XRB), formed via accretion onto a compact object from a companion, exhibit a wide array of variability on timescales ranging from milliseconds to days to months. In particular, most XRBs are transient, meaning that they alternate between intervals of quiescence and outburst due to changes in the rate of mass accretion. For example, in Be X-ray binaries consisting of a Be star and an accreting neutron star, these periods of outburst are caused by transient accretion when the compact object passes near its companion in its orbit. XRBs hosting black holes undergo weeks- or months-long outbursts due to slow accumulation of material into their accretion disks, leading to a critical point at which material accretes much more efficiently onto the compact object (Mineshige and Wheeler, 1989).
@@
These accretion-powered outbursts are distinct from bursting behavior caused by thermonuclear burning. Neutron star XRBs which persistently accrete from low mass companions at a relatively low level may accumulate material on their surfaces without burning until a critical depth is reached, producing runaway burning of accreted hydrogen and helium. We observe this runaway conflagration in the form of a rapid flash known as a Type I X-ray burst, lasting only seconds to minutes, during which the source luminosity may increase by multiple orders of magnitude up to the Eddington luminosity. The depth at which accumulated material ignites plays a decisive role in the duration and structure of the resulting burst. Some sources exhibit “superbursts” which resemble Type I X-ray bursts but which last several hours and decay on the timescale of days due to ignition at much larger column depths producing extended carbon burning (in’t Zand, 2017; Cumming, 2003; Keek et al., 2008; Serino et al., 2016). Due to their extended timescale, superbursts can provide an opportunity to probe how accretion disks and coronae respond to high levels of X-ray irradiation. The rarity of these events, however, has thus far limited investigations into superbursts. In this paper, we present an analysis of the newly-discovered source, MAXI J1752-457, which was observed by the Nuclear Spectroscopic Telescope Array (NuSTAR) during the decay phase of a superburst.
@@
The first detection of MAXI J1752-457 with the Monitor of All-sky X-ray Image (MAXI) Gas Slit Camera (GSC) occurred during the scan transit at 18:23 on November 9, 2024 UTC (MJD 60623.766) with a 4–10 keV flux of $982\pm 50$ mCrab, or $(1.12\pm 0.06)\times 10^{-8}$$\mathrm{erg\,cm^{-2}\,s^{-1}}$(Serino et al., 2024). The flux dropped rapidly over subsequent scan transits. Several months prior to the detection of MAXI J1752-457 with MAXI, an X-ray transient, deemed EP240809a, was detected by Einstein Probe (EP) with a position consistent with MAXI J1752-457 (Liu et al., 2024). This earlier period of brightening, which was observed on August 9, 2024 (UTC), reached a much lower flux than the November burst: Liu et al. (2024) reported an unabsorbed 0.4–5 keV flux of $5.3\times 10^{-11}$$\mathrm{erg\,cm^{-2}\,s^{-1}}$, corresponding to about 1.5 mCrab. An optical flare from the source was also observed by the Asteroid Terrestrial-impact Last Alert System (ATLAS) beginning about three weeks before the EP detection (Zhu et al., 2024). The authors pointed out that the soft X-ray and the blue optical spectra were consistent with a tidal disruption event (TDE), but no host galaxy was found. Therefore, the nature of the source remained unknown.
@@
MAXI/GSC observations showed a thermal spectrum which softened during the rapid decay, leading Negoro et al. (2024) to propose that the origin of the source activity was a long X-ray burst from an uncatalogued X-ray binary, rather than a TDE. By combining MAXI/GSC and NinjaSat observations of the November burst, Aoyama et al. (2025) have shown that the source is indeed consistent with an accreting neutron star. The flux evolution of the burst, which exhibits a fast rise and exponential decay on the timescale of days, along with an X-ray spectrum which can be described with a spherical blackbody, indicate that the transient event corresponded to accretion-induced thermonuclear burning. In particular, the hours-long timescale of the burst and the years-long recurrence time (inferred from the lack of previous detections of the source by MAXI), indicate a superburst.
@@
At the time that it was detected by MAXI/GSC, MAXI J1752-457 had a small Solar Aspect Angle (SAA) of $45^{\circ}$, making follow-up observations impossible for most pointing X-ray observatories. Thanks to the ability of NuSTAR to point near or at the Sun, we were able to trigger hard X-ray observations of the MAXI confidence region. These observations confirmed that MAXI J1752-457 is coincident with EP240809a, and preliminary spectral analysis provided the first evidence for a hard X-ray excess in addition to the blackbody spectral component (Pike et al., 2024).
@@
Here, we present a detailed analysis of two NuSTAR observations of MAXI J1752-457 following its November, 2024, superburst, and we place the flux of the source as measured by NuSTAR in the context of MAXI/GSC monitoring of the burst. In Section II, we briefly present the MAXI/GSC observations of the burst. In Section III, we present both spectral and timing analyses of the source with NuSTAR. In Section IV, we discuss the implications of these results, and finally we list our conclusions in Section V. All errors are quoted at the $90\%$ confidence level unless otherwise stated.
II MAXI light-curve
@@
MAXI, launched in 2009, is an all-sky X-ray monitor installed on the Japanese Experiment Module onboard the International Space Station (Matsuoka et al., 2009). It consists of two instruments: the Gas Slit Camera (GSC) with a net bandpass of 2–20 keV and field of view (FOV) of $1.5^{\circ}\times 160^{\circ}$ (Mihara et al., 2011), and the Solid-state Slit Camera (SSC) with a net bandpass of 0.7–7 keV and FOV of $1.5^{\circ}\times 90^{\circ}$. The SSC ceased operations on June 22, 2022, and was therefore not used during observations of MAXI J1752-457. MAXI/GSC scans the full X-ray sky once every orbit (about 92 minutes), such that individual sources are within the FOV for about $45$ seconds per orbit (Sugizaki et al., 2011). The resulting 3-$\sigma$ detection limit for one scan is $80$ mCrab in the 3–10 keV band (Negoro et al., 2016). The MAXI/GSC light-curve of MAXI J1752-457 available from the online MAXI light-curve portal [^1] http://maxi.riken.jp/star_data/J1752-457/J1752-457.html contained data with a low signal-to-noise ratio due to contamination by the bright low-mass XRB (LMXB), H 1735-444, separated from MAXI J1752-457 by $2.8^{\circ}$ in the GSC anode direction. Therefore, we instead obtained the MAXI/GSC light-curve using the point-spread-function fit method presented by Morii et al. (2016). We excluded scanning data if none of the cameras completely covered the source region. We show the MAXI/GSC light-curve for MJD 60622 – MJD 60629 in the top panel of Figure 1. We have re-binned the data by up to 10 orbits for clarity before and after the brightest phase of the burst. We note that visual inspection of the MAXI/GSC light-curve of MAXI J1752-457 binned into 16 days over the course of more than 15 years reveals a hint of a weak enhancement ($\sim$ 1–2 mCrab) beginning in July, 2024, and lasting roughly through November, 2024.
III NuSTAR Data and Analysis
@@
MAXI J1752-457 was observed twice by NuSTAR during the tail of the 2024 superburst. The first observation was performed using a tiling strategy designed to locate the source within the MAXI/GSC 90% confidence region. Four tiling observations were carried out with exposures of about $5$ ks each. The source was detected in one of these observations with OBSID 81010302001. Using the resulting localization, a second, pointed observation (OBSID 81001347001) was carried out less than a day later. The start and end times and the total exposure times of the two NuSTAR observations are listed in Table 1. We extracted cleaned event lists for both observations using the standard nupipeline options with saamode=NONE and tentacle=no. We used HEASOFT v6.34, NuSTARDAS v2.1.4a, and CALDB v20241104.
@@
During both NuSTAR observations, the source was at a low SAA, leading to complications in the image reconstruction process. At low SAA, the primary star tracker, Camera Head Unit 4 (CHU4), which is located on the X-ray optics bench, may be blinded by the Sun and therefore become unusable for image reconstruction. The usual science mode, which utilizes CHU4, is called Mode 1. Outside of Mode 1, image reconstruction must be performed using combinations of CHUs 1, 2, and 3, which are located on the spacecraft bus. This is referred to as Mode 06 data. Depending on the pointing of the spacecraft any given time, one of seven combinations of these three star trackers may be used to produce an attitude solution. Due to thermal flexing of the mountings of CHUs 1, 2, and 3, the resulting solution may include inaccuracies which manifest as a “smearing” of the source or the presence of multiple centroids in the reconstructed image. In order to most effectively extract scientific products from Mode 06 data, we used the nusplitsc command to split the cleaned event list according to active CHU combination. We performed nusplitsc two separate times, once with splitmode=NORMAL and another time with splitmode=STRICT. In some cases, strict splitting can help to remove flickering between CHU combinations, therefore mitigating multiple centroids. However, based on visual inspection using SAOImageDS9 (Joye and Mandel, 2003), we determined that using the strict splitmode did not improve the image reconstruction compared to normal mode, and in fact strict mode only resulted in a reduction in the overall number of events available for science products. We therefore proceeded using the event files produced using splitmode=NORMAL.
@@
We used DS9 to inspect the event files from Mode 1 and the split event files from Mode 06. For each file, we chose a circular source extraction region and centered it on the source using the DS9 centroid calculation feature. During the first observation, there was significant stray light which did not overlap with the source position, but which was close enough to the source to become a consideration when choosing the source region. In order to avoid stray light contamination of the source data, we chose source regions with radius $35^{\prime\prime}$ for FPMA data and $40^{\prime\prime}$ for FPMB data from the first observation. For both FPMA and FPMB data from the second observation, we chose source regions with radius $45^{\prime\prime}$. For all event files, we chose circular background regions with radius $100^{\prime\prime}$. The stray light region, characteristic centroid smearing, and examples of the source and background regions we chose for FPMA are shown in the Mode 06 images for both observations in Figure 2. After selecting extraction regions, we used nuproducts to produce source and background spectra and light-curves for each cleaned event file. The background-subtracted light-curves (summed FPMA and FPMB) are shown in the bottom panels of Figure 1. Finally, for each observation, we used addspec to sum the Mode 1 and Mode 06 data into one pair of FPMA source and background spectra and one pair of FPMB spectra. We used ftgrouppha to rebin the spectra according to the method described by Kaastra and Bleeker (2016) with a minimum of 10 counts per bin.
III.1 Spectral Analysis
@@
Table 2: Spectral parameters and unabsorbed 2–10 keV flux measured during the NuSTAR observations. footnotetext: Radius calculated assuming a distance of 88 kpc, without correcting for scattering in the neutron star atmosphere.
@@
[…]
$f_{\mathrm{2-10}}\,\mathrm{(erg\,cm^{-2}\,s^{-1})}$ $7^{+1}_{-2}\times 10^{-11}$ $4.3^{+0.3}_{-0.4}\times 10^{-11}$
@@
Given the superburst interpretation presented in Aoyama et al. (2025), we added a physically-motivated single-temperature blackbody component to our spectral models, representing thermal emission from the surface of a neutron star heated by either thermonuclear burning during the burst or by persistent accretion flow. Notably, the peak flux observed by MAXI/GSC during the burst was around $2\times 10^{-8}$$\mathrm{erg\,cm^{-2}\,s^{-1}}$. Although the distance to MAXI J1752-457 is unknown, the source is located at $(\alpha,\delta)=(268.269,-45.866)$, or (l,b)=(346.092,9.8517)(l,b)=(346.092,-9.8517), placing it at a $17^{\circ}$ separation from the Galactic center. Like Aoyama et al. (2025), we tentatively assume that the angular proximity of the source to the Galactic center translates into physical proximity, giving a distance of around 8 kpc. This assumption leads to a derived peak luminosity on the order of $10^{38}$$\mathrm{erg\,s^{-1}}$, very near the Eddington luminosity for spherical accretion onto a neutron star: $L_{\mathrm{Edd}}=1.38\left(M/{M_{\odot}}\right)\times 10^{38}$$\mathrm{erg\,s^{-1}}$, where the canonical neutron star mass is $M=1.4M_{\odot}$. Thus, the physical interpretation of the source as a bursting neutron star is consistent with a distance of around 8 kpc, and we therefore present derived luminosities and blackbody radii using this fiducial value.
@@
We also considered slightly different forms of this basic spectral model. For example, we replaced the single-temperature blackbody with a multi-temperature disk blackbody. This model resulted in a comparable or slightly worse test statistic (e.g. $\chi^{2}/\mathrm{d.o.f.}=112/91$ for the second observation). This model yields an inner disk temperature of around $kT_{\mathrm{in}}=0.7$ keV and a slightly shallower power law with a photon index of $\Gamma=3.4$. Similarly, replacing the simple power law component with a thermally Comptonized continuum component (nthcomp; Zdziarski et al., 1996; Życki et al., 1999), did not improve the test statistic ($\chi^{2}/\mathrm{d.o.f.}=110/91$ for the second observation). In this case, the relevant spectral parameters did not change compared to the simple power law model.
In order to check the self-consistency of the spectral model, we also reproduced the spectral fits using simpl (Steiner et al., 2009) rather than an analytic power law. We again found no appreciable change in the test statistic ($\chi^{2}/\mathrm{d.o.f.}=110/91$ for the second observation). This model yields a slightly larger blackbody radius (at 8 kpc, uncorrected for atmospheric scattering effects) of $11^{+2}_{-1}$ km and $7.4^{+0.6}_{-0.5}$ km for the first and second NuSTAR observations, respectively. For both observations, using simpl instead of an analytic power law results in a decrease in the blackbody temperature by a few percent, an increase in the photon index of about 0.1, and an upscattering fraction of around $10\%$, assuming that the power law is seeded by the spherical blackbody component.